OpenStudy (anonymous):

CHECK ANSWER An investor wants to analyze the earnings of a mutual fund account. Three years ago, the value of the account was $36,000 and it is now worth $50,400. If the account is compared to a bank account paying interest that is compounded continuously, what interest rate rounded to the nearest hundredth of a percent would the bank account have to pay to match the mutual fund account’s earnings? (Assume the only deposit was to open the account.) R=14.346?

OpenStudy (anonymous):

Continuously compounded means that the interest for an account is being calculated and applied constantly. The basic formula for continuously compounded interest is A=P*e^(r*t). Where P is the principle invested, e is the exponential function, r is the interest rate, t is the time period of investment, and A is the total amount after time t. So, in your question, 50400 = 36000*e^(3r) e^(3r) = 50400/36000 3r = ln(50400/36000) r = (ln(50400/36000))/3 Which works out to about 11.2%, so your R wouldn't work.